OpenStudy (anonymous):

Find the angle if cos = 4/7 and sin is negative

OpenStudy (anonymous):

Angle is between 270 and 360. Cos^-1 of 4/7 = 55.15 Angle is 360-55.15=304.85 degrees

OpenStudy (anonymous):

sin=sqrt(33)/7

OpenStudy (anonymous):

and if u punch in the calculator for sin^(-1) you'll find the angle